OpenStudy (anonymous):

If we roll a fair six-sided die three times, and let X be the number of results that are divisible by 2, what is the probability that X=2?

OpenStudy (perl):

well this is a big sample space

OpenStudy (anonymous):

yes!

OpenStudy (perl):

for three die, X can equal 0 , 1 , 2, or 3 . X = 0 means all three dice are not divisible by 2 X = 1 means 1 of the dice is divisible by 2 X = 2 means 2 of the dice are divisible by 2 X = 3 means all three dice are divisible by 2

OpenStudy (anonymous):

riight okay gotcha

OpenStudy (perl):

but we want to do 3 dice. ok so lets see

OpenStudy (anonymous):

well we are rolling 1 dice three times

OpenStudy (perl):

divisible by 2 means 'even' , so lets use even. for X = 2 we have (even, even , odd) or ( even , odd, even) or (odd, even , even)

OpenStudy (perl):

now we count up for (even, even, odd ) we have 3 x 3 x 3 for (even, odd, even) we have 3 x 3 x 3 for ( odd, even, even) we have 3 x 3 x 3 . now there are a total of 6x6x6 possible ways to throw three die. so ( 3 x 3 x 3 x 3 ) / ( 6 x 6 x 6 )

OpenStudy (perl):

this is because there are 3 even numbers 2,4,6 and 3 odd numbers 1,3,5

OpenStudy (anonymous):

okay, i see!

OpenStudy (perl):

assuming i read the question correctly, it could be more clearly phrased

OpenStudy (anonymous):

thank you for your help. i appreciate it!

OpenStudy (perl):

:D

OpenStudy (zarkon):

you can also see it as a binomial dist...3 trials...success probability 1/2 thus \[P(X=2)= {3 \choose 2} \left(\frac{1}{2}\right)^3=\frac{3}{8}\]

OpenStudy (perl):

yes that works too , that is probably easier than my approach

OpenStudy (perl):

hmm, this is because each throw is independent and probability of even is 1/2

OpenStudy (anonymous):

binomial dist is what unit i am in now. Thanks @zarkon